Marie has 39 quarters and nickels. The total value is $ 5.15.
Write an equation and solve for how many quarters and nickels she has.

Answer by ikleyn(52781)

.
Marie has 39 quarters and nickels. The total value is $ 5.15.
Write an equation and solve for how many quarters and nickels she has.
~~~~~~~~~~~~~~~~~~~~~

Let Q be the number of quarters.
Then the number of nickels is (39-Q).


Write the total money equation, in cents

    25Q + 5*(39-Q) = 515  cents.


Simplify and find Q

    25Q + 5*39 - 5Q = 515

    25Q - 5Q = 515 - 5*39

      20Q    =    320

        Q    =    320/20 = 16.


ANSWER.  16 quarters and 39-16 = 23 nickels.


CHECK.  25*16 + 5*23 = 515 cents, in total.    ! correct !

Solved.

---------------------

For introductory coin problems,  look into the lessons
    - Coin problems
    - More Coin problems
    - Solving coin problems without using equations
    - Kevin and Randy Muise have a jar containing coins
in this site.

Read them attentively and become an expert in this field.